Having trouble with getting breakdown on this, can anyone help?
Facing rising fax costs, a firm issued guidelines that transmissions of 10 pages or more should be done by 2 day mail instead. Exceptions are allowed, they want the average to be 10 or below. A firm examined 35 randomly chosen fax transmissions during the next year, yielding a sample mean of 14.44 with a standard deviation of 4.45 pages. (a) At the .01 level of significance, is the true mean greater than 10? (b) Use excel to find the right tail – p value.




Answer by stanbon(75887)

Facing rising fax costs, a firm issued guidelines that transmissions of 10 pages or more should be done by 2 day mail instead. Exceptions are allowed, they want the average to be 10 or below. A firm examined randomly chosen fax transmissions during the next year, yielding a sample mean of 14.44 with a standard deviation of 4.45 pages. (a) At the .01 level of significance, is the true mean greater than 10? (b) Use excel to find the right tail – p value.
---------------
Ho: mu = 10
Ma: mu > 10
---------------------
one-tailed test with alpha=1%; critical value = 2.326
-------------
Test Statistic:
t(14.44) = (14.44-10)/(4.45/sqrt(35)) = 4.44*sqrt(35)/4.45 = 5.90
------------
The test statistic is greater than the critical value so reject Ho.
-------
The p-value = 0.0000000018236...
====================================
Cheers,
Stan H.